OpenStudy (anonymous):

What is the first step in solving the equation three-tenths equals x divided by three plus one-tenth for x? Add one-tenth to both sides of the equal sign. Add three-tenths to both sides of the equal sign. Subtract one-tenth from both sides of the equal sign. Subtract three-tenths from both sides of the equal sign.

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (math&ing001):

3/10 = x/3 + 1/10 3/10 - 1/10 = x/3 2/10 = x/3 So option 4

OpenStudy (math&ing001):

option 3*

OpenStudy (anonymous):

3/10 = x/3 + 1/10 --- subtract 1/10 from both sides you want to get the x term by itself